4/5 of 5 kg apples were used on Monday. The next day 1/3 of what was left was used. Weight (in kg) of apples left now is (a) 2/7 (b) 1/14 (c) 2/3 (d) 4/21

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (c) is correct here
From the question it is given that,
4/5 of 5 kg apples were used on Monday = (4/5) × 5

= 20/5
= 4 kg

Then,
The next day 1/3 of what was left was used = (1/3) × 1

= 1/3 kg
So, Weight (in kg) of apples left now is = 1–(1/3)
= (3 - 1)/3
= 2/3 kg of apples


Thanks